OpenStudy (anonymous):

how do you simplify (4^8)^3 ?

OpenStudy (anonymous):

@DavidUsa @Etevian @Lady.Liv1776 @DeadShot

OpenStudy (zarkon):

\[=4^84^84^8=\cdots\]

OpenStudy (anonymous):

THANK U

OpenStudy (owlcoffee):

\[(C^a)^b=C ^{ab}\]

OpenStudy (davidusa):

Exactly. So 4^32

OpenStudy (zarkon):

\[8\cdot 3=24\] \[(4^8)^3=4^{24}\]